For the circuit shown, find the magnitude of the potential difference between the points a and b. The resistances R1 and R2 are 102.0 ? and 135.5 ? respectively. The batteries are both held at voltage V = 0.5000 V2 2 1 1
Answer #1

Batteries are in series

=> circuit voltage = 2 * 0.5 = 1 V

Resistors are all in series

=> circuit resitance = 2 * (102 + 135.5) = 475 ohm

=> circuit current = 1/475 = 0.0021 A

potential difference between the points a and b

Vab = 2*R2*current + 0.5 V

= 2*135.5*0.0021 + 0.5 V

= 1.07 volt.